Reactive Power Generation has the following capital structure. Its corporate tax rate is 21%.

Security Market Value Required Rate
of Return
Debt $ 20 million 2 %
Preferred stock 30 million 4
Common stock 50 million 8

What is its WACC? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Total value=(20+30+50)=$100 million

After-tax rate of debt=2*(1-tax rate)

=2*(1-0.21)=1.58%

WACC=Respective return*Respective weight

=(20/100*1.58)+(30/100*4)+(50/100*8)

=5.52%(Approx).